Prev Next

SW&IT Meeting Minutes 2020-04-27

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200420
LR20200420 - Bec, Matthieu 40.0
LR20200420 - Holmberg, Carl 32.0
LR20200420 - Edmeades, David 40.0
LR20200420 - Sola, Igor 40.0
LR20200420 - Hooper, Stephen 80.0
LR20200420 - Zhang, Yang 40.0
LR20200420 - Zhang, Xianyu 0.0

News from management

  • Christian's daily 10am vs. Pat's mountain coordination

Previous Action Items

Owner Action Status
IS/DE create a new Luci simulator for ROWG purpose (+ work out resources available on the VM) existing OSA simulator will be fine
DE/XZ Xianyu's IDL personal license WIP - DE with Louise
MB SOUL/UAO follow-up DONE - shared presentation and discussed git with Arcetri

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • board meeting
    • meeting ALTA, Arcetri SOUL in git, AGW theta binary encoder
    • DMS bug fixed ALTA
    • AO in git branches renaming
  • This week
    • ROWG#6 planning and meeting schedule
    • phone interviews, review ad for senior
    • cobbler / kickstart, ims f30
    • DMS
      • weather loss, websockets, AdSec, etc.
      • LBC stiched png preview
    • ARGOS meeting

XZ

  • Last week
    • AO SW versions merge plan and discussion
    • check and dicuss AO SW backwards compatiable with LN//ARGOS
    • PB action items startup and discuss with MG
    • Meetings
  • This week
    • PB action items
    • Merge UAO to SOUL
    • PB SW/HW

DE

  • Sent request for task to Pat re: LeRoy working on Luci servers, sent LeRoy list of images that I want
  • Double-check me on time estimate for LeRoy doing the upgrade work during SSD (I initially estimated 1 day)
  • Xianyu IDL license
    • Very expired
    • To upgrade, easiest/cheapest is for new license
    • Sent details
  • Remote setup of new vibration analysis hardware for PG
  • Mod osa-luci
Next:
  • Veeam agent quotes for Ovirt
  • IDL purchase?
  • Investigate state of educational Autodesk licenses in prep for major licensure change they are pushing
  • Work on new vibration laptop
  • Robert VPN issues--I think he's connecting to the wrong server but getting that info out is challenging. I'll remote in if it goes on much longer.
  • JH MS Office license (self-stated low priority, waiting on him)
  • KVM troubleshooting
  • Work out VM migration to VMHost1

YZ

  • Previous:
    • Continue IT#8021, LBC rotator trailing issues. Not much activities and progress. Checked some TCS/LBC code. Olga and John Hill added some more notes.
    • Finished a new TCS/IIF interface document with SHARK command info added. Now waiting on Robert Reynolds to make it official.
    • AO-Cravat work: setup github repo for all to download and collaborate, also working on an update to the Python Strehl code.
    • modsAlign setup on home machine, did some testing and experiment with DS9 XPAget. Current conclusion: we can get rid of pyraf (use xpaget and Python subprocess), but modsAlign needs more work to become Python3 ready.
    • TMS Mathematica -> Python conversion (Andrew replied, needs some time to get back with more information).
    • Sent a summary of open MODS WCS keyword problems to Olga.
    • Finished several safety training courses.
  • This week:
    • See what more can be done with IT#8021.
    • Continue working on AO-Cravat Python code update.
    • Maybe follow up with Al and Steve see how do they think about the current code?
    • Continue working on modsAlign (Python3 updating and xpaget refactoring).
    • Follow up on TMS Mathematica questions.
    • Follow up on MODS WCS problems with Olga.
    • Hoping to finish some more safety training.

SH

  • mountain power down and power up
  • Upgraded rm507 hosts to Fedora 31
  • Installed new computer for rm507 monitor testing
  • Add cabling for Tucson VPN
  • meetings

next:

  • ROWG monitor
  • move docker to ovirt cluster
  • web1.tucson, abell.as to ovirt cluster
  • TCS Fedora 31 build/test machine (and that frees up TCS-test)
    • TCS test to mountain with Fedora 31 install
    • then upgrade tcs2 to Fedora 31
    • then upgrade tcs1 to Fedora 31
  • DNS to ovirt cluster

CH

  • Last week:
    • AGW: meeting with Bec, Bucki, Cavallaro, Cox, Haddad, & Morris regarding way forward on replacing obsolete UMAC encoder boards.
      • Options: order one new board to play with now & search for used encoder boards, dive in and order a full set of replacement boards.
      • There will be several months lag time between new board order and delivery.
      • In the meantime, J. Morris will search on-line for more of the older grey encoder boards.
      • Use the lab M3 AGW to test software and wiring configuration.
    • DMS: implemented repurposed sensor data storage, added newly installed sensors.
    • indi-core: completed software testing.
  • This week:
    • ARGOS: fix rpm packaging.
    • so-nightwriter: investigate webifying.

IS

Last week:
  • PIT
    • INAF PIT 2.1004.0 test release finally send to INAF for testing
    • Worked on PIT admin tool (authorization)
  • OT
    • MODS twilights issues on scrip generation.
    • LUCI lamps xml element apparently creating problems in LUCI simulator. More test on that this week.
  • LUCI
    • GEIRS software installation in two of the luci simulators, osa-luci.sim and isola-luci.
    • Going through the GEIRS code to get familiar with it.

This week:
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • work with Andrew on current test OT release issues
  • LUCI
    • keep digging a little bit more on the latest GEIRS code.
    • Look into LUCI WebIO device configuration.
Topic revision: r9 - 27 Apr 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback